OpenStudy (anonymous):

f(x,y)=(x/(y^2)) + xy whats the partial derivatives for this function?

OpenStudy (amistre64):

derive with respect to x; then derive with respect to y ...

OpenStudy (amistre64):

when working with respect to x; consider y to be a constant; and same concept for with respect to y

OpenStudy (anonymous):

i got this but im not sure its correct.. fx(x,y) = (y^2/ y^4 ) + y and fy(x,y)=-1/2(y^2)^(-1/2) . 2y + x

OpenStudy (amistre64):

\[f(x,y)=xy^{-2} + xy\] \[f_x(x,y)=y^{-2} + y\] \[f_y(x,y)=-2xy^{-3} +x\] if i see it right

OpenStudy (amistre64):

treat the "other" variables as tho they were just a constant

OpenStudy (anonymous):

thank you ! i skipped the function "transformation" to xy^-2..

OpenStudy (anonymous):

i cant find the good answer button and ive tried 3 different browsers lol what am i doing wrong